LASER SPECTROSCOPY OF A HALOCARBOCATION: OBSERVATION OF THE ELECTRONIC SPEC-TRUM OF CH_{2}I^{+}

\underline{S. A. REID}, C. TAO, AND C. MUKARAKATE, Department of Chemistry, Marquette University, Milwaukee, WI 53233.

We report the first observation of the gas-phase electronic spectrum of a simple halocarbocation, CH_{2}I^{+}. Fluorescence excitation and emission spectra were measured under jet-cooled conditions using a pulsed discharge source with CH_{2}I_{2} as precursor. The spectral carrier was identified by: (1) comparison of experimentally derived ground state vibrational frequencies for various isotopomers with the predictions of Density Functional Theory (DFT) calculations, and (2) simulation of the observed rotational contour in our excitation spectra, which indicated an a-type transition, consistent with an excited state of A_{1} symmetry. Calculations predict a triplet A_{1} excited state with transition energy in the visible region, and the observed transition is therefore assigned as ^{3}A_{1}-^{1}A_{1}, with band origin at 15180 cm^{−1}.